There are three industrial firms in Happy Valley.

Firm

Initial Pollution Level

Cost of Reducing Pollution by 1 Unit

(Units)

(Dollars)

A 30 20
B 40 30
C 20 10

The government wants to reduce pollution by 60 units, so it gives each firm 20 tradable pollution permits.

Because firmB   has the highest cost of reducing pollution by 1 unit, it would like tobuy 20 permits from   another firm. Because firmC   has the lowest cost of reducing pollution by 1 unit, it is willing tosell 20 permits to   the firm with the highest cost.

Complete the second column of the following table by indicating who buys and sells permits in this case. Complete the fourth column of the table with the number of permits each firm ends up with after trading with the others. Next, subtract the number of permits from the initial pollution level to determine how many units of pollution reduction each firm has to complete, and enter these values into the fifth column of the table. Next, determine the total cost of reducing the pollution levels you determined in the previous column, and enter that into the final column of the table. Finally, add up the total pollution reduction costs for each firm and enter the total amount in the final cell of the total row in the following table.

Firm

Buys or Sells

Initial Pollution Level

Permits after Trade

Remaining Pollution Reduction

Cost of Reducing Pollution by 1 Unit

Total Pollution Reduction Cost

(Dollars)

(Dollars)

A 30 20
B    40 30
C    20 10
Total

If permits could not be traded, firm A would have to remove

units of pollution at a total cost of

, firm B would have to remove

units of pollution at a total cost of

, and firm C would have to remove

units of pollution at a total cost of

. The overall cost of reducing pollution in this case is

, which is   than with tradable permits.
Answer #1

Because firm B has the highest cost of reducing pollution by 1 unit, it would like to buy 20 permits from another firm. Because firm C has the lowest cost of reducing pollution by 1 unit, it is willing to sell 20 permits to the firm with the highest cost.

Firm

Buys or Sells

Initial Pollution Level

Permits after Trade

Remaining Pollution Reduction

Cost of Reducing Pollution by 1 Unit

Total Pollution Reduction Cost

(Dollars)

(Dollars)

A Neither buy nor sell 30 20 30-20 =10 20 10*20 = 200
B Buy 40 40 40-40 =0 30 0
C Sell 20 0 20-0 =20 10 20*10 =200
Total 400

If permits could not be traded, firm A would have to remove 10 units of pollution at a total cost of 10*20 = 200, firm B would have to remove 20 units of pollution at a total cost of 20*30 = 600, and firm C would have to remove 0 units of pollution at a total cost of 0. The overall cost of reducing pollution in this case is 800, which is Greater than with tradable permits.